# {cutout}02.753cm0cm1 empty ViC-Bench     : Benchmarking Visual-Interleaved Chain-of-Thought Capability in MLLMs with Free-Style Intermediate State Representations

### v2

@@ 3.4 Embodied Long-Horizon Planning
Stage 1. We regard the final visual state achieved after completing the sequence as the correct option. Three incorrect options are generated using the ambiguous IVS, which are extracted from the original trace and manually refined to increase difficulty, as displayed in Fig. 2 (c). This setup thoroughly tests the long-horizon planning capabilities of models, providing baselines for latter evaluations.
@@
Stage 2. For each sample, we take the initial observation of action sequence in Stage 1 as the starting state $\mathbf{O}_{s}$, and the completely achieved state as the ending state $\mathbf{O}_{e}$. MLLMs in this stage are required to open-endedly respond with the action sequence which can successfully transition from $\mathbf{O}_{s}$ to $\mathbf{O}_{e}$.
Stage 3 & IVS Generation. Based on the construction priors in Stages 1 & 2, we employ human experts to perform Stage 3 assessment in the real-world scenarios. In this way, we deploy the human-machine-in-the-loop paradigm, where MLLM as a planner and human as an actioner [82], to conduct free-style IVS enhanced evaluations. Concretely, human experts capture the intermediate images based on the model’s thinking plans and upload them to the cloud storage via a dedicated mobile APP, subsequently utilizing function calls for the step-wise collaborative reasoning evaluation.
3.5 Complex Counting
@@
Raw Data & Pre-processing. We build this task based on the samples from JHU-CROWD++ [57] and CoMT [13]. According to annotations in [13], we first filter out samples with headcounts ranging from 10 to 40, then remove those that are ambiguous or difficult to delineate. As displayed in Fig. 2 (d), we then manually draw irregular lines to separate the input image into four regions labeled as $\mathbf{\{1,2,3,4\}}$, leading to the completed input image $\mathbf{H}$. In the end, we totally construct 250 samples.
@@
Stage 1. We manually count the visible heads and regard the correct headcounts as the ground truth option, tackling limitations in previous work [13] that solely relies on DNNs for annotations. We then employ human experts to establish three incorrect options, leading to higher selection difficulties.
@@
Stages 2 & 3. The input format of image remains consistent with Stage 1, but the QA format changes from multiple-choice to open-ended, where we require models to directly respond with the headcounts for each region. The output format is defined as <Begin> ($\mathbf{1}$, $\mathbf{h_{1}}$), ($\mathbf{2}$, $\mathbf{h_{2}}$), ($\mathbf{3}$, $\mathbf{h_{3}}$), ($\mathbf{4}$, $\mathbf{h_{4}}$) </End>, where $\mathbf{h_{i}}$ ($\mathbf{i}\in\{1,2,3,4\}$) denotes the headcount in each separated region.
@@
IVS Generation. Considering that complex counting does not hold the explicit dynamics of action sequences like maze navigation or jigsaw puzzle, and the empirical experimental results further indicate that overly fine-grained masks can lead to overthinking and object hallucinations, we thus utilize the coarse-grained region-aware mask instead of fine-grained single head mask for the IVS generation. As a result, we manually draw bounding boxes for $\mathbf{H}$ to accurately cover masks over all the heads in each separated region, consistently leading to four IVS for each data sample.
@@
Figure 3: The overall illustration of our evaluation workflow taking maze navigation as the example.
4 Evaluation Suite
4.1 Progressive Three-Stage Evaluation Strategy
@@
As illustrated in Fig. 3, we propose a holistic progressive three-stage evaluation strategy, rather than simply focusing on the final answer, to examine the untamed understanding of VI-CoT capability in MLLMs. Stage 1 only focus on the final results, employing the common multiple-choices QA [53, 18, 42] to preliminarily determine the visual CoT performance. Based on POMDP [82], Stage 1 can be formulated as:
@@
where $p_{sys}$ and $p_{task}$ denote system and task prompts. $p_{cot}$ and $p_{io}$ refer to CoT and IO prompts, designating the inner thought flow and output format. $(Q,\mathbf{v_{0}},C)$ represents question and inital observation with options. $\mathbf{r_{i}}$ ($\mathbf{i}\in\{1,...,\mathbf{T}\}$) are the reasoning rationales. $\mathbf{C_{S}}$ is the selection. In Stage 2, we convert the multiple-choices QA into the more challenging open-ended format, i.e.,
@@
where $\mathbf{A_{F}}$ is the formatted answer. Doing so allows us to observe the reasoning performance in an open-ended manner, which leads to direct subjective Answer-Only evaluations, thereby serving as an important indicator for Stage 3. Finally, Stage 3 pays attention to the legal free-style evaluations using function calls in the agent-form to explore in-depth thinking gains brought by the IVS representations, addressing the shortcomings in previous works where fixed IVS might forcefully influence the inherent planning in MLLMs by constraining the judgment path. The overall process can be represented as:
@@
$\pi_{\theta}(p_{sys},(Q,\mathbf{v_{0}}),((\mathbf{a_{1}},\mathbf{v_{1}}),...,(\mathbf{a_{t}},\mathbf{v_{t}})),p_{task},p_{cot},p_{io})\rightarrow(\mathbf{a_{t+1}},\mathbf{v_{t+1}}),$ (5)
where $\mathbf{R}$ denotes the final answer and $(\mathbf{a_{t}},\mathbf{v_{t}})$ refers to action and IVS feedback at the $t^{th}$ step. Our free-style exploration in the agent-form further stimulates the impact of IVS, thereby more comprehensively investigating the untamed VI-CoT capability inherented in current MLLMs.
4.2 Evaluation Metrics
@@
Stage 1. We mainly focus on whether the final choice is correct or not, consistently utilizing ACC as the sole metric across tasks. As shown in Fig. 3 (a), we first perform rule-based extraction to obtain the selected options, then directly compare them with Ground Truth to compute the ACC metric, i.e.,
@@
where $TP$, $TN$, $FP$, and $FN$ denote the number of true positives, true negatives, false positives, and false negatives. We can tentatively observe reasoning performance, offering extensive references.
@@
[…]
@@
In maze navigation, we extract the final predictions by the rule-based method, then determine the legal endpoint $\mathbf{P_{f}}$ reached by the predicted path using our rendering simulations. We then compare $\mathbf{P_{f}}$ with Ground Truth since each constructed maze has a unique shortest path. If $\mathbf{P_{f}}$ exists in the coordinate set $\mathbf{C_{set}}$ of Ground Truth, we thus determine $\mathcal{S}_{\text{reached}}^{k}$ based on the distance from $\mathbf{P_{f}}$ to the starting point $\mathbf{S}$. If $\mathbf{P_{f}}$ is not in $\mathbf{C_{set}}$, we directly take $\mathbf{R\_o}$ as 0. In jigsaw puzzle, we follow maze navigation to extract final predictions. Considering the uniqueness of puzzle correspondence sequence, we directly perform global-aware strict matching under the component-wise paradigm between predictions and Ground Truth. We calculate $\mathbf{R\_o}$ through taking the number of correctly filled patches as $\mathcal{S}_{\text{reached}}^{k}$ and the counts of extracted patches as $\mathcal{S}^{k}$. Considering the atomic action of embodied long-horizon planning task have dependency constraints, we verify the correctness of atomic action utilizing semantic matching [14] rather than strict matching, keeping consistent with [83]. A pair-wise match is regarded as successful only when the semantic similarity of atomic action pair exceeds 0.95. For ACC, the sample is taken as correct if all the atomic actions are validly matched. For $\mathbf{R\_o}$, we take the accumulated effective length of verified actions and the total predicted steps as $\mathcal{S}_{\text{reached}}^{k}$ and $\mathcal{S}^{k}$, respectively. As for complex counting, we follow jigsaw puzzle to perform global-aware pair-wise matching, except that we extra implement a fault tolerance mechanism due to the high difficulty of this task for current MLLMs, with a tolerance set to 1.
The compliance with explicit rules in maze navigation and jigsaw puzzle poses challenges to the instruction-following and visual perception capabilities of MLLMs, which correlatively promotes VI-CoT evaluation, leading to our Legality metric. In maze navigation, we consider two types of illegal behaviors: going out of bounds and hitting walls. For jigsaw puzzles, the illegal behaviors are: repeated patch placements and repeated filling of vacant regions. Specifically, we segment the predictions followed by conducting step-wise simulation rendering to determine the legal steps, i.e.,
@@
Stage 3. Building upon Stages 1 & 2, we further introduce a new metric denoted ThinkGain to examine the influence of free-style IVS on VI-CoT reasoning performance. Drawing inspirations from the rewards of GRPO in [24], we black-box the retrospection process and focus only on the decision states ($\mathbf{D_{d}}$ $\mathbf{D_{o}}$) both before the retrospection commences and after it concludes, avoiding the negative impact of numerous ongoing factors on VI-CoT evaluation. We then employ the Recall metric defined in Stage 2 to assess $\mathbf{D_{d}}$ and $\mathbf{D_{o}}$. Overall, the ThinkGain metric can be represented as:
@@
where $\mathbf{R\_d}$ denotes the Recall metric calculated with $\mathbf{D_{d}}$. Besides, the definition of $\mathbf{D_{d}}$ varies due to inconsistence in task representations. In maze navigation, we regard $\mathbf{D_{d}}$ as the terminal point reached before the first retrospection. In jigsaw puzzle and complex counting, we treat the state of each patch or region upon its first utilization as $\mathbf{D_{d}}$. For embodied long-horizon planning, we directly employ human experts to determine two decision states.
4.3 Incremental Prompting Information Injection Strategy
Based on three-stage evaluations and metrics, we introduce the Incremental Prompting Information Injection (IPII) strategy to ablatively explore the prompting factors for VI-CoT under three varying levels of global-aware prompting information. Concretely, Level-1 establishes a baseline only using the original prompts of three-stage evaluations. Level-2 involves implicit VI-CoT prompts, which guides models to update its internal step-wise IVS, thereby compelling models to leverage its visual imagination for subsequent planning trajectory, rather than solely relying on the initial observation. Level-3 further augments prompts with external knowledge, providing supporting information to externally enhance the visual perception.
5 Experiments
Table 2: The performace evaluations of advanced MLLMs in our metrics (%) on maze navigation. Note that we highlight the best performance in bold and underline the second performance.
@@ 5 Experiments
GPT-4o-20240513 [47] OpenAI $\times$ 50.00 0.00 11.65 53.18 66.40 69.41 33.01 63.76
o3 [48] OpenAI 87.60 17.52 29.11 70.55 74.40 76.86 44.05 81.33
o4-mini [49] OpenAI 94.40 15.45 22.33 68.45 58.40 61.78 27.98 69.88
Gemini-2.5-pro-preview-03-25 [63] Google 94.00 6.80 12.95 55.05 68.80 70.27 32.96 73.28
Gemini-2.0-flash [62] Google 60.80 0.00 11.35 54.94 53.20 59.48 28.11 64.79
Claude-3.5-Sonnet-20241022 [3] Anthropic 53.20 1.60 11.80 52.07 25.60 27.90 6.23 84.63
Step-1o-vision-32k [58] StepFun $\times$ 46.40 0.80 10.72 51.08 16.40 21.64 4.84 30.87
Doubao-1.5-vision-pro-32k [7] ByteDance $\times$ 44.00 1.20 10.89 52.03 13.20 28.53 6.22 58.40
SenseChat-Vision-V5.5 [54] SenseTime $\times$ 62.80 0.00 11.22 53.03 22.40 34.36 10.50 63.92
Hunyuan-Vision-20250103 [68] Tencent $\times$ 30.00 0.40 8.89 48.73 16.40 27.86 7.03 46.57
@@
LLaVA-OneVision-72B [32] ByteDance & NTU $\times$ 38.00 0.00 8.77 52.25 18.40 30.44 8.17 49.80
InternVL-2.5-78B [12] Shanghai AI Lab $\times$ 46.80 0.00 10.75 53.47 29.60 36.31 11.27 60.52
InternVL-3-78B [87] Shanghai AI Lab $\times$ 43.60 0.00 11.74 52.73 37.60 45.03 12.21 61.86
Qwen2-VL-72B-Instruct [70] Alibaba $\times$ 47.20 0.00 13.67 52.94 45.20 50.33 8.11 53.58
Qwen2.5-VL-72B-Instruct [5] Alibaba $\times$ 55.60 1.20 11.27 49.13 41.20 45.30 11.77 57.04
Qwen-Omni-Turbo [67] Alibaba $\times$ 28.00 0.00 7.97 30.89 14.00 22.84 5.52 43.44
GLM-4V-Plus [64] Zhipu AI $\times$ 54.40 0.00 10.53 51.57 32.80 37.46 8.08 50.38
SkyWork-R1V1-38B [50] Skywork 36.40 0.40 7.91 38.49 3.60 12.55 0.27 29.34
Table 3: The performace evaluations of advanced MLLMs on jigsaw puzzle.

### v3

@@ III-D Embodied Long-Horizon Planning
Stage 1. We directly regard the final visual state achieved after completing the overall action sequence as the correct option. Three incorrect options are generated using ambiguous IVS, which are extracted from the original execution trace and manually refined to increase perceptual and semantic difficulty, as displayed in Fig. 2 (c). All options are further verified by human experts to ensure plausibility. This setup thoroughly evaluates the long-horizon planning capabilities of models and provides strong baselines for subsequent evaluations.
@@
Stage 2. For each sample, we take the initial observation of the action sequence from Stage 1 as the starting state $\mathbf{O}_{s}$, and the fully achieved final state as the ending state $\mathbf{O}_{e}$. In this stage, MLLMs are required to open-endedly generate an action sequence that successfully transitions from $\mathbf{O}_{s}$ to $\mathbf{O}_{e}$.
Stage 3 & IVS Generation. In this stage, we leverage the rendering engine [ai2thor] to enable dynamic evaluation. Specifically, each action predicted by models is rendered and executed within the environment. An action is deemed legal only if it is successfully executed, thereby strictly penalizing hallucinated or physically infeasible operations. Complementing this quantitative evaluation, we further introduce a human-in-the-loop paradigm for qualitative assessment in real-world settings. Following [magebench], we adopt a collaborative framework in which the MLLM serves as the planner and the human acts as the executor. Within this setup, human experts capture intermediate visual states (IVS) based on the model’s responses and upload them to the cloud via a dedicated mobile APP. These IVS are then utilized in conjunction with function calling to enable step-wise, human-machine collaborative evaluation.
III-E Complex Counting
@@
Raw Data & Pre-processing. We build this subset based on the samples from JHU-CROWD++ [Jhu-crowd++] and CoMT [Comt_AAAI]. According to the data annotations in [Comt_AAAI], we first filter out samples with headcounts ranging from 10 to 40, then remove those that are ambiguous or difficult to delineate. As displayed in Fig. 2 (d), we then manually draw irregular lines to separate the input image into four regions, leading to the completed input image $\mathbf{H}$. In the end, we construct 250 diverse samples.
@@
Stage 1. We manually count the visible heads and regard the correct headcounts as the ground truth option, tackling limitations in previous work [Comt_AAAI] that solely relies on DNNs for annotations. We then employ human experts to establish three incorrect options, leading to higher selection difficulties.
@@
Stages 2 & 3. The input image remains consistent with Stage 1, but the QA format changes from multiple-choice to open-ended, where we require models to directly respond with the headcounts for each region. The output format is defined as <Begin> ($\mathbf{1}$, $\mathbf{h_{1}}$), ($\mathbf{2}$, $\mathbf{h_{2}}$), ($\mathbf{3}$, $\mathbf{h_{3}}$), ($\mathbf{4}$, $\mathbf{h_{4}}$) </End>, where $\mathbf{h_{i}}$ ($\mathbf{i}\in\{1,2,3,4\}$) denotes the headcount in each region.
@@
IVS Generation. Considering that complex counting lacks the explicit action dynamics present in tasks like maze navigation or jigsaw puzzle, and empirical results further indicate that overly fine-grained masks can induce overthinking and object hallucinations, we thus utilize the coarse-grained region-aware masks instead of fine-grained per-head masks for IVS generation. Specifically, for each region, we manually draw a number of bounding boxes according to the model’s predicted head count, regardless of whether this matches the ground-truth number. This process yields sequentially stacked masks that can fully cover, under-cover, or over-cover the actual heads, thereby reflecting the model’s counting behavior.
@@
Figure 3: The overall illustration of our progressive three-stage evaluation strategy taking maze navigation as the example.
IV Evaluation Suite
IV-A Progressive Three-Stage Evaluation Strategy
@@
As illustrated in Fig. 3, we propose a holistic progressive three-stage evaluation strategy, rather than simply focusing on the final answer, to examine the untamed understanding of VI-CoT capability in MLLMs. Remarkably, we take the maze navigation task as an example to demonstrate the key steps in the entire workflow. Stage 1 only focuses on the final results, employing the common multiple-choices QA [Scienceqa, ding2024mmvqa, 76_masry2022chartqa] to preliminarily determine the visual CoT performance. Based on POMDP [magebench], Stage 1 can be formulated as:
@@
where $p_{sys}$ and $p_{task}$ denote system and task prompts. $p_{cot}$ and $p_{io}$ refer to CoT and IO prompts, designating the inner thought flow and output format. $(Q,\mathbf{v_{0}},C)$ represents question and initial observation with options. $\mathbf{r_{i}}$ ($\mathbf{i}\in\{1,...,\mathbf{T}\}$) are the reasoning rationales. $\mathbf{C_{S}}$ is the selection. In Stage 2, we convert the multiple-choices QA into the more challenging open-ended format, i.e.,
@@
where $\mathbf{A_{F}}$ is the formatted answer. Doing so allows us to observe the reasoning performance in an open-ended manner, which leads to direct subjective Answer-Only evaluations, thereby serving as an important indicator for Stage 3. Finally, Stage 3 pays attention to the legal free-style evaluations using function calls in the agent-form to explore in-depth thinking gains brought by the IVS representations, addressing the shortcomings in previous works where fixed IVS might forcefully influence the inherent planning in MLLMs by constraining the judgment path. Overall, this process can be represented as:
@@
$\begin{split}\pi_{\theta}\big(&p_{sys},(Q,\mathbf{v_{0}}),((\mathbf{a_{1}},\mathbf{v_{1}}),\dots,(\mathbf{a_{t}},\mathbf{v_{t}})),\\ &p_{task},p_{cot},p_{io}\big)\rightarrow(\mathbf{a_{t+1}},\mathbf{v_{t+1}}),\end{split}$ (5)
where $\mathbf{R}$ denotes the final answer and $(\mathbf{a_{t}},\mathbf{v_{t}})$ refers to action and IVS feedback at the $t^{th}$ step. Our free-style exploration in the agent-form further stimulates the influence of IVS, thereby more comprehensively investigating the untamed VI-CoT capability inherented in advanced MLLMs.
IV-B Evaluation Metrics
@@
Stage 1. We mainly focus on whether the final choice is correct or not, consistently utilizing ACC as the sole metric across four tasks. As displayed in Fig. 3 (a), we first perform rule-based extraction to obtain the selected options, then directly compare them with Ground Truth to compute the ACC metric, i.e.,
@@
where $TP$, TNTN, FPFP, and $FN$ denote the number of true positives, true negatives, false positives, and false negatives. In this way, we can tentatively observe reasoning performance, offering extensive references.
@@
$k_{0}$ $=\underset{k}{\arg\max}\frac{\left|\mathcal{S}_{\text{reached}}^{k}\right|}{\left|\mathcal{S}^{k}\right|},$ (7)
$\mathbf{R\_o}$ $=\frac{\left|\mathcal{S}_{\text{reached}}^{k_{0}}\right|}{\left|\mathcal{S}^{k_{0}}\right|},$ (8)
@@
In maze navigation, we extract the final predictions by the rule-based method, then determine the legal endpoint $\mathbf{P_{f}}$ reached by the predicted path using our rendering simulations. We then compare $\mathbf{P_{f}}$ with Ground Truth since each constructed maze has a unique shortest path. If $\mathbf{P_{f}}$ exists in the coordinate set $\mathbf{C_{set}}$ of Ground Truth, we thus determine $\mathcal{S}_{\text{reached}}^{k}$ based on the distance from $\mathbf{P_{f}}$ to the starting point $\mathbf{S}$. If $\mathbf{P_{f}}$ is not in $\mathbf{C_{set}}$, we directly take $\mathbf{R\_o}$ as 0. In jigsaw puzzle, we follow maze navigation to extract final predictions. Considering the uniqueness of puzzle correspondence sequence, we directly perform global-aware strict matching under the component-wise paradigm between predictions and Ground Truth. We calculate $\mathbf{R\_o}$ through taking the number of correctly filled patches as $\mathcal{S}_{\text{reached}}^{k}$ and the counts of extracted patches as $\mathcal{S}^{k}$. Considering the atomic action of embodied long-horizon planning task have dependency constraints, we verify the correctness of atomic action utilizing semantic matching [paddlenlp] rather than strict matching, keeping consistent with [Embodied-reasoner]. Notably, non-critical steps (i.e., observe and move forward) are ignored during matching. A pair-wise match is regarded as successful only when the semantic similarity of atomic action pair exceeds 0.95. For ACC, the sample is taken as correct if all the atomic actions are validly matched. For $\mathbf{R\_o}$, we take the accumulated effective length of verified actions and the total predicted steps as $\mathcal{S}_{\text{reached}}^{k}$ and $\mathcal{S}^{k}$, respectively. As for complex counting, we follow jigsaw puzzle to perform global-aware pair-wise matching, except that we further implement an explicit fault-tolerance mechanism due to the high difficulty of this task for current MLLMs, with a tolerance threshold set to 1.
The compliance with explicit rules in maze navigation, jigsaw puzzle, and embodied long-horizon planning tasks poses challenges to the instruction-following and visual perception capabilities of MLLMs, which correlatively promotes VI-CoT evaluation, leading to our Legality metric. In maze navigation, we consider two types of illegal behaviors, namely going out of bounds and hitting walls. For jigsaw puzzles, illegal behaviors include repeated patch placements and repeated filling of vacant regions. Regarding embodied long-horizon planning, legal steps necessitate valid actions performed on interactable objects. Specifically, we segment the predictions followed by conducting step-wise simulation rendering to determine the legal steps, i.e.,
@@
Stage 3. Building upon Stages 1 & 2, we further introduce a new metric denoted ThinkGain to examine the influence of free-style IVS on VI-CoT performance. Drawing inspirations from the reward system of GRPO in [Deepseek-r1], we black-box the retrospection process and focus only on the decision states ($\mathbf{D_{d}}$ & $\mathbf{D_{o}}$) both before the retrospection commences and after it concludes, avoiding the negative impact of numerous ongoing factors on VI-CoT evaluation. We then employ the Recall metric defined in Stage 2 to assess $\mathbf{D_{d}}$ and $\mathbf{D_{o}}$. Overall, the ThinkGain metric can be represented as:
@@
where $\mathbf{R\_d}$ denotes the Recall metric calculated with $\mathbf{D_{d}}$. Besides, the definition of $\mathbf{D_{d}}$ varies due to inconsistence in task representations. In maze navigation, we regard $\mathbf{D_{d}}$ as the terminal point reached before the first retrospection. In jigsaw puzzle and complex counting, we treat the state of each patch or region upon its first utilization as $\mathbf{D_{d}}$. For embodied long-horizon planning, we define $\mathbf{D_{d}}$ by identifying reflective adjustments in the execution path, specifically capturing repetitive actions such as re-navigating to the same location or re-picking up the same object.
IV-C Incremental Prompting Information Injection Strategy
Based on above three-stage evaluations and metrics, we introduce the Incremental Prompting Information Injection (IPII) strategy, formally represented as a set of hierarchical prompting levels $\mathcal{H}=\{\mathcal{P}_{\text{L1}},\mathcal{P}_{\text{L2}},\mathcal{P}_{\text{L3}}\}$, to ablatively explore the prompting factors for VI-CoT performance. Concretely, Level-1 establishes the baseline only utilizing the original instruction set $\mathcal{I}_{base}$, defined as $\mathcal{P}_{\text{L1}}=\mathcal{I}_{base}$. Level-2 involves implicit VI-CoT prompts by injecting a guidance term $\mathcal{I}_{imp}$, formulated as $\mathcal{P}_{\text{L2}}=\mathcal{P}_{\text{L1}}\oplus\mathcal{I}_{imp}$. This is deployed to guide the models to update its internal step-wise IVS, thereby compelling them to leverage visual imagination for subsequent planning trajectory rather than solely relying on the initial visual observation. Level-3 further augments the prompts with external knowledge $\mathcal{K}_{ext}$ to enhance visual perception, which can be represented as $\mathcal{P}_{\text{L3}}=\mathcal{P}_{\text{L2}}\oplus\mathcal{K}_{ext}$.
V Experiments
TABLE II: The performace evaluations of advanced MLLMs in terms of targeted metrics (%) on maze navigation. Note that we highlight the best performance in bold and underline the second performance.
@@ V Experiments
GPT-4o-20240513 [openai2024gpt4o] OpenAI $\times$ 50.00 0.00 11.65 53.18 66.40 69.41 33.01 63.76
o3 [openai2025o3mini] OpenAI 87.60 17.52 29.11 70.55 74.40 76.86 44.05 81.33
o4-mini [openai2025o4mini] OpenAI 94.40 15.45 22.33 68.45 58.40 61.78 27.98 69.88
Gemini-2.5-pro-preview-03-25 [Gemini-2.5-pro-preview-03-25] Google 94.00 6.80 12.95 55.05 68.80 70.27 32.96 73.28
Gemini-2.0-flash [Gemini-2.0-flash] Google 60.80 0.00 11.35 54.94 53.20 59.48 28.11 64.79
Claude-3.5-Sonnet-20241022 [anthropic2024claude3d5] Anthropic 53.20 1.60 11.80 52.07 25.60 27.90 6.23 84.63
Step-1o-vision-32k [step-1o-vision-32k] StepFun $\times$ 46.40 0.80 10.72 51.08 16.40 21.64 4.84 30.87
Doubao-1.5-vision-pro-32k [Doubao-1.5-vision-pro-32k] ByteDance $\times$ 44.00 1.20 10.89 52.03 13.20 28.53 6.22 58.40
SenseChat-Vision-V5.5 [SenseChat-Vision-V5.5] SenseTime $\times$ 62.80 0.00 11.22 53.03 22.40 34.36 10.50 63.92
Hunyuan-Vision-20250103 [Hunyuan-Vision-20250103] Tencent $\times$ 30.00 0.40 8.89 48.73 16.40 27.86 7.03 46.57
@@
LLaVA-OneVision-72B [Llava-onevision] ByteDance & NTU $\times$ 38.00 0.00 8.77 52.25 18.40 30.44 8.17 49.80
InternVL-2.5-78B [Internvl] Shanghai AI Lab $\times$ 46.80 0.00 10.75 53.47 29.60 36.31 11.27 60.52
InternVL-3-78B [InternVL3] Shanghai AI Lab $\times$ 43.60 0.00 11.74 52.73 37.60 45.03 12.21 61.86
Qwen2-VL-72B-Instruct [Qwen2-vl] Alibaba $\times$ 47.20 0.00 13.67 52.94 45.20 50.33 8.11 53.58
Qwen2.5-VL-72B-Instruct [Qwen2d5-vl] Alibaba $\times$ 55.60 1.20 11.27 49.13 41.20 45.30 11.77 57.04
Qwen-Omni-Turbo [Qwen-Omni-Turbo] Alibaba $\times$ 28.00 0.00 7.97 30.89 14.00 22.84 5.52 43.44
GLM-4V-Plus [glm2024chatglm] Zhipu AI $\times$ 54.40 0.00 10.53 51.57 32.80 37.46 8.08 50.38
SkyWork-R1V1-38B [Skywork-r1v] Skywork 36.40 0.40 7.91 38.49 3.60 12.55 0.27 29.34
V-A Experimental Setup
Evaluation Models. We globally select a total of 18 top-performing MLLMs for comprehensive evaluation, comprising 10 commercial proprietary models and 8 powerful open-source models. Regarding proprietary models, we include the leading OpenAI and Gemini series, specifically GPT-4o-20240513 [openai2024gpt4o], o3 [openai2025o3mini], o4-mini [openai2025o4mini], Gemini-2.5-pro [Gemini-2.5-pro-preview-03-25], and Gemini-2.0-flash [Gemini-2.0-flash]. We also assess popular models from other competitive organizations, including Claude-3.5-Sonnet [anthropic2024claude3d5], Step-1o-vision-32k [step-1o-vision-32k], Doubao-1.5-vision-pro-32k [Doubao-1.5-vision-pro-32k], SenseChat-Vision-V5.5 [SenseChat-Vision-V5.5], as well as Hunyuan-Vision-20250103 [Hunyuan-Vision-20250103]. As for open-source models, we select representative series with their largest parameter capacities to investigate performance gaps. These include LLaVA-OneVision-72B [Llava-onevision], InternVL-2.5-78B [Internvl], InternVL-3-78B [InternVL3], Qwen2-VL-72B-Instruct [Qwen2-vl], Qwen2.5-VL-72B-Instruct [Qwen2d5-vl], Qwen-Omni-Turbo [Qwen-Omni-Turbo], and GLM-4V-Plus [glm2024chatglm]. Finally, we evaluate MLLMs with targeted thinking capabilities, represented by SkyWork-R1V1-38B [Skywork-r1v]. Note that we take GPT-4o [openai2024gpt4o] as the baseline model for our experiments.
Implementation Details. We access proprietary and open models via APIs and local deployments. The maximum token limit is 8192, temperature is 0, both Top-K and Top-P are 1. The rest of hyper-parameter settings keep same with the default settings of VLMEvalKit [Vlmevalkit]. All the experiments are conducted on a machine with 8 $\times$ NVIDIA A100 GPUs (80G). Due to space constraint, we only apply IPII strategy for maze navigation here. Specifically, Level-1 prompts are consistent with origin prompts, which are illustrated in Fig. 1 of supplementary material. Level-2 builds upon Level-1 by incorporating the following prompting information: Please make sure that after executing the move at each step, you should update your internal intermediate visual state, rather than remaining in the initial input visual state, as displayed in Fig. 5 of supplementary material. As shown in Fig. 6 of supplementary material, Level-3 further builds upon Level-2 by explicitly incorporating the coordinates information of the starting and end points in maze navigation.
TABLE III: The performace evaluations of advanced MLLMs on jigsaw puzzle.